### v1

@@ III-B2 Automated Network Initialization
Furthermore, NetQMPI handles the synchronization of the quantum control flow. While the raw SDK requires manual flush() invocations, NetQMPI primitives include implicit flushing mechanisms or offer explicit flush() operations to ensure state consistency across the network.
III-C Point to point functions
The fundamental requirement of a distributed quantum network is the ability to transfer quantum states between nodes. NetQMPI encapsulates the standard Quantum Teleportation protocol into two atomic primitives: qsend and qrecv.
@@ III-C2 The qrecv Primitive
This encapsulation drastically simplifies the development process. As shown in the example below, the complex coordination required by the raw SDK (discussed in Section II) is reduced to intuitive, single-line commands:
@@ 4 ROOT_RANK = 0
18
19 m = local_qs[0].measure()
20 conn.flush()
21
22 print(f"Rank {rank}: measured {m}")
@@ III-D2 Expose and Unexpose
To address this, NetQMPI introduces the expose and unexpose primitives, as proposed in the architectural framework cited [Cardama2026NetQIR:Computing]. Instead of copying the state, the expose operation creates a shared context where a specific qubit (or set of qubits) becomes conceptually accessible to other processes for collective operations, without violating the no-cloning principle.
@@ IV Comparative and Use Cases
IV-A Benchmark Definition
The goal is to generate the maximally entangled state $|\text{GHZ}\rangle=\frac{1}{\sqrt{2}}(|00\dots 0\rangle+|11\dots 1\rangle)$ shared among $N$ distinct network nodes.
@@ IV-A Benchmark Definition
    One root node (e.g., Rank 0) prepares a qubit in superposition ($|+\rangle$).
@@
IV-B Qualitative Analysis: Programming Complexity
@@
1. NetQMPI (This Work): Following the MPI standard, the solution is written in a single source file. The logic scales automatically with the number of processes. The developer uses the rank variable to distinguish between the root (Control) and the others (Targets). The distributed CNOT is abstracted, or the expose primitive is used to create the shared state in a single call. Complexity: $O(1)$ in terms of file management. The code remains constant regardless of $N$.
2. NetQASM SDK (SquidASM / SimulaQron): As these tools operate on a role-based paradigm, the developer must explicitly define the behavior for each node. For a 3-node GHZ state, this requires writing three separate application flows (e.g., app_alice.py, app_bob.py, app_charlie.py). Furthermore, a Remote CNOT is not a native primitive. The developer must manually implement the ”telegate” protocol: generating EPR pairs, performing local measurements, sending classical bits via a separate socket, and applying corrections. Complexity: $O(N)$ in terms of source files and code maintenance. Extending the system from 3 to 4 nodes requires writing a completely new script for the new node and updating the root script to handle the new connection.
@@
IV-C Code Implementation Examples
@@
IV-C1 NetQMPI (SPMD Paradigm)
@@
IV-C2 NetQASM SDK (Role-Based)
@@
Crucially, note that the developer must explicitly instantiate the connection objects (EPRSocket for quantum links and Socket for classical communication) before the main logic. This reveals the scalability bottleneck: adding a third node (Charlie) forced us to:
@@
IV-C3 NetSquid (Discrete-Event / Hardware)
NetSquid operates at a lower abstraction level. Before implementing the GHZ protocol, the user must define the physical hardware. Listing 3 illustrates the verbosity of setting up just the nodes and channels. The actual algorithm requires defining a class inheriting from NodeProtocol for each entity.
@@ IV-C3 NetSquid (Discrete-Event / Hardware)
IV-C4 QuNetSim (High-Level Simulator)
@@
IV-D Quantitative Analysis: lines of code (LOC)
@@
Framework Files
LOC
LOC
Paradigm
@@
Implications for Quantum Software Engineering: This reduction in code complexity is critical for the development of distributed quantum applications. By decoupling the program logic from the network size, NetQMPI allows for rapid prototyping and ensures a bug-free implementation of complex protocols. In a domain where debugging distributed quantum states is notoriously difficult, minimizing the codebase surface reduces the potential for implementation errors, allowing researchers to focus on algorithmic correctness rather than infrastructure management.
V Conclusions
@@
  • Paradigm SPMD: NetQMPI facilitates a transition from the labor-intensive, role-based programming model of existing SDKs to a Single Program Multiple Data (SPMD) paradigm. By unifying the control logic into a single source file where roles are defined by the process rank, we have demonstrated that the complexity of setting up a quantum network drops from $\mathcal{O}(N^{2})$ to $\mathcal{O}(1)$ in LOC number. This decoupling of program logic from network size is a fundamental requirement for scalable quantum software engineering.
@@
  • Integration with the quantum ecosystem: A critical objective of NetQMPI is its seamless integration with the existing quantum stack. By compiling down to the standardized NetQASM ISA, NetQMPI acts as a backend-agnostic middleware. This allows researchers to execute high-level code on state-of-the-art simulators like SquidASM and NetSquid, leveraging their rigorous physical noise models without facing their steep learning curves. Furthermore, this architecture ensures that applications written today in NetQMPI will be executable on future physical quantum hardware that supports the NetQASM standard.
In summary, NetQMPI represents a significant step forward in quantum software engineering. It allows developers to write readable, maintainable, and bug-free distributed quantum applications and paving the way for the implementation of complex distributed algorithms.
@@
[Uncaptioned image] Tomás F. Pena (Senior Member) received the Ph.D. degree in physics from the University of Santiago de Compostela (USC), Santiago, Spain, in 1994. He is currently an Full Professor with the Department of Electronics and Computer Science, USC, and a Member of the Research Center in Intelligent Technologies (CiTIUS), Santiago de Compostela, Spain. His research interests include distributed quantum computing, high-performance computing, parallel systems architecture, optimization of performance in irregular codes and with sparse matrices, and Big Data technologies.

### v2

@@ IV-B2 Automated Network Initialization
Furthermore, NetQMPI handles the synchronization of the quantum control flow. While the NetQASM SDK requires manual flush() invocations, NetQMPI primitives include implicit flushing mechanisms or offer explicit flush() operations to ensure state consistency across the network.
IV-C Point to point functions
The fundamental requirement of a distributed quantum system is the ability to transfer quantum states between nodes. NetQMPI encapsulates the standard Quantum Teleportation Protocol into two atomic primitives: qsend and qrecv.
@@ IV-C2 The qrecv Primitive
This encapsulation drastically simplifies the development process. As shown in Figure 8, the complex coordination required by the raw SDK (discussed in Section III) is reduced to intuitive, single-line commands.
@@ 4 ROOT_RANK = 0
16
17 m = local_qs[0].measure())
@@ IV-D2 Expose and Unexpose
To address this, NetQMPI introduces the expose and unexpose primitives, as proposed in [Cardama2026NetQIR:Computing]. Instead of copying the state, the expose operation creates a shared context where a specific qubit (or set of qubits) becomes conceptually accessible to other processes for collective operations, without violating the no-cloning principle.
@@ V Comparative and Use Cases
V-A Benchmark Definition
The goal is to generate the maximally entangled state $|\text{GHZ}\rangle=\frac{1}{\sqrt{2}}(|00\dots 0\rangle+|11\dots 1\rangle)$ shared among $N$ distinct network nodes, where each rank holds exactly one qubit.
@@ V-A Benchmark Definition
    One root node (e.g., Rank 0) prepares one qubit in superposition ($|+\rangle$).
@@
V-B Qualitative Analysis: Programming Complexity
@@
1. NetQMPI (This Work): Following the SPMD paradigm, the solution is written in a single source file. The logic scales automatically with the number of processes. The developer uses the rank variable to distinguish between the root (Control) and the others (Targets). The distributed CNOT is abstracted and the expose primitive is used to create the shared state in a single call. Complexity: $O(1)$ in terms of file management. The code remains constant regardless of $N$.
2. NetQASM SDK (SquidASM / SimulaQron): As these tools operate on a role-based paradigm, the developer must explicitly define the behavior for each node. For a 3-node GHZ state, this requires writing three separate application flows (e.g., app_alice.py, app_bob.py, app_charlie.py). Furthermore, a Remote CNOT is not a native primitive. The developer must manually implement the ”telegate” protocol, which involves generating EPR pairs, performing local measurements, sending classical bits via a separate socket, and applying corrections. Complexity: $O(N)$ in terms of source files and code maintenance. Extending the system from 3 to 4 nodes requires writing a completely new script for the new node and updating the root script to handle the new connection.
@@
V-C Code Implementation Examples
@@
V-C1 NetQMPI (SPMD Paradigm)
@@
V-C2 NetQASM SDK (Role-Based)
@@
Crucially, note that the developer must explicitly instantiate the connection objects (EPRSocket for quantum links and Socket for classical communication) before the main logic, which reveals the scalability bottleneck: adding a fourth node (Dave) forced us to:
@@
V-C3 NetSquid (Discrete-Event / Hardware)
NetSquid operates at a lower level of abstraction. Before implementing the GHZ protocol, the user must define the physical hardware. Listing 3 illustrates the verbosity of setting up just the nodes and channels. The actual algorithm requires defining a class inheriting from NodeProtocol for each entity.
Due to the excessive length and technical complexity of the full implementation, the remaining algorithmic code is omitted. It is crucial to emphasize that this comparison does not aim to undermine the utility of discrete-event simulators like NetSquid, but rather to highlight the conceptual gap between physical layer modeling and high-level application development. These tools serve fundamentally different purposes: while discrete-event engines are indispensable for validating hardware fidelity and noise models, they are not designed for algorithmic prototyping. Thus, the level of detail presented here demonstrates that these are not entities directly comparable to NetQMPI, but instead underlying layers that our proposed approach aims to abstract away.
@@ V-C3 NetSquid (Discrete-Event / Hardware)
V-C4 QuNetSim (High-Level Simulator)
@@
V-D Quantitative Analysis: lines of code (LOC)
@@
Framework Files LOC (Network) LOC (Comp.) Paradigm
@@
Implications for Quantum Software Engineering: This reduction in code complexity is critical for the development of distributed quantum applications. By decoupling the program logic from the network size, NetQMPI allows for rapid prototyping and ensures a bug-free implementation of complex protocols. In a domain where debugging distributed quantum states is notoriously tricky, minimizing the codebase surface reduces the potential for implementation errors, allowing researchers to focus on algorithmic correctness rather than infrastructure management.
VI Conclusions
@@
  • SPMD paradigm: NetQMPI facilitates a transition from the labor-intensive, role-based programming model of existing SDKs to a Single Program Multiple Data (SPMD) paradigm. By unifying the control logic into a single source file where roles are defined by the process rank, we have demonstrated that the complexity of setting up a quantum network drops from $\mathcal{O}(N^{2})$ to $\mathcal{O}(1)$ in LOC number. This decoupling of program logic from network size is a fundamental requirement for scalable quantum software engineering.
@@
  • Operational Implementation of QMPI: While previous proposal for a QMPI remain theoretical specifications, NetQMPI provides a functional software implementation. This work demonstrates the practical viability of the MPI paradigm for quantum networks, moving beyond abstract definitions to provide a working execution environment. Crucially, the implementation resolves the conflict between classical broadcasting semantics and the No-Cloning Theorem, proving that high-level distributed abstractions can be effectively mapped to rigorous physical instructions via the NetQASM ecosystem.
  • Integration with the quantum ecosystem: A critical objective of NetQMPI is its seamless integration with the existing quantum stack. By compiling down to the standardized NetQASM ISA, NetQMPI acts as a backend-agnostic middleware. This allows researchers to execute high-level code on state-of-the-art simulators, such as SquidASM and NetSquid, leveraging their rigorous physical noise models without facing their steep learning curves. Furthermore, this architecture ensures that applications written today in NetQMPI will be executable on future physical quantum hardware that supports the NetQASM standard.
In summary, NetQMPI represents a significant step forward in quantum software engineering. It enables developers to write readable, maintainable, and bug-free distributed quantum applications, paving the way for the implementation of complex distributed algorithms.
@@
[Uncaptioned image] Tomás F. Pena (Senior Member) received the Ph.D. in Physics from the University of Santiago de Compostela (USC), Santiago, Spain, in 1994. He is currently Full Professor with the Department of Electronics and Computer Science, USC, and a Senior Member of the Research Center in Intelligent Technologies (CiTIUS), Santiago de Compostela, Spain. His research interests include distributed quantum computing, high-performance computing, parallel systems architecture, optimization of performance in irregular codes and with sparse matrices, and Big Data technologies.